368070:

Marguerite wants to rent a carpet cleaner. Company A rents a carpet cleaner for $15 per day. Company B rents a carpet cleaner for $100 per week, with a one-time fee of $5. The following functions represent the rate structures of the two rental companies: x = the number of weeks Company A f(x) = 15(7x) Company B g(x) = 100x + 5 The function h(x) = f(x) – g(x) represents the difference between the two rate structures. Determine which statements about h(x) and about renting a carpet cleaner are true. Check all that apply. If Marguerite rents for 2 weeks, it will cost her more if she rents from Company B. If Marguerite rents for 2 weeks, it will cost her less if she rents from Company B. If Marguerite rents for 1 week, it will cost her the same at either company. If Marguerite rents for 1 week, it will cost her more if she rents from Company A. h(x) = 5x – 5 h(x) = 5x + 5

Extrinix:

So, let's assume that we can just plug in 2 weeks, so, f(x)=15(7(2)) = $210 g(x)=100(2)+5 = $205 Those are for 2 weeks, now for 1 week f(x)=15(7(1)) = $105 g(x)=100(1)+5 = $105 Notice something? Company A has a $5 increase from Company B So you can choose, `If Marguerite rents for 2 weeks, it will cost her less if she rents from Company B.` You can also choose, `If Marguerite rents for 1 week, it will cost her the same at either company.` Now for this last one, just plug in your variables, h(x)=15(7(2)) - 100(2)+5 = $15 difference Now plug it into your options, h(x)=5(2)-5 = 5 h(x)=5(2)+5 = 15 so, your options that are chosen would be, `If Maguerite rents for 2 weeks, it will cost her less if she rents from Company B.` `If Maguerite rents for 1 week, it will cost her the same at either company.` `h(x)=5x+5`

368070:

Thanks bro